# Leveraging Atom Loss Errors in Fault Tolerant Quantum Algorithms

### v2

@@ Abstract
Errors associated with qubit loss constitute an important source of noise in many quantum hardware systems, particularly in neutral atom quantum computers. We develop a theoretical framework to handle these errors in logical algorithms, incorporating decoding techniques and circuit-level optimizations. Focusing on experimentally-motivated error models, we introduce a delayed-erasure decoder which leverages information from state-selective readout to accurately correct loss errors, even when their precise locations are unknown. Our decoding technique is compatible with a wide range of quantum error correction codes and general logical circuits. Using this decoder, we identify strategies for detecting and correcting atom loss based on the logical circuit structure. For deep circuits with a large number of gate layers prior to logical measurements, we explore methods to integrate loss detection into syndrome extraction with minimal overhead, identifying optimal strategies depending on the qubit loss fraction in the noise. In contrast, many algorithmic subroutines involve frequent gate teleportation, shortening the circuit depth before logical measurement and naturally replacing qubits without additional overhead. We simulate such a teleportation-based algorithm, involving a toy model for small-angle synthesis and find a significant improvement in logical error rates as the loss fraction increases, with loss handled solely through teleportation. These results provide a path forward for advancing large-scale fault tolerant quantum computation in systems with loss errors.
@@ 1 Introduction
Quantum error correction (QEC) is believed to be essential for realizing large-scale quantum computation, as it enables suppression of errors [1, 2, 3]. However, its practical implementation remains challenging due to its substantial resource overhead. Recent experimental advancements have demonstrated remarkable progress in implementing quantum operations across multiple logical qubits [4, 5, 6, 7, 8], and operating below error thresholds [9]. These advances make it clear that practical QEC performance can be substantially improved by tailoring the error correction strategy to the particular experimental error model [10, 9, 11], choice of logic gates [4], and the structure of the algorithm itself [12, 13].
@@
In particular, qubit loss and leakage from computational subspace are dominant noise sources in many hardware systems, and can have a dramatic effect on the practical performance of quantum processors [14, 15]. In the absence of loss correction, all qubits within a QEC code can eventually disappear, destroying the encoded quantum information. Conversely, directly detecting loss provides valuable information about the error location, in contrast with Pauli errors, which are indirectly inferred from syndrome information. Recent work has shown that the rich information provided by so-called erasure detection can in fact improve QEC performance substantially [16, 17, 18, 19, 10, 11, 20, 21, 22, 23, 24, 25, 26]. Similarly, noise bias present in many platforms [14, 4, 9] can also be leveraged to improve performance [27, 28, 29, 30]. These studies motivate developing strategies that incorporate loss detection and harness bias, to improve practical QEC performance.
@@
In this Article, we explore the role of loss and biased Pauli errors in error-corrected circuits, with a focus on neutral atom quantum computers. We leverage loss detection upon measurement and design QEC techniques to benefit from loss with minimal experimental overhead. In particular, since the detection and correction of erasure is delayed for several gate operations, we develop a delayed-erasure decoder to accurately interpret logical measurement results from the measured syndromes and loss detections, despite uncertainty in the exact location of the loss error.
@@
Using this delayed-erasure decoder, we investigate the impact of loss errors in logical circuits. We demonstrate that the algorithmic structure significantly influences the optimal strategy to detect and correct loss, as summarized in Fig. 1(a). Concretely, for high-depth circuits prior to logical measurement, we develop hardware-efficient methods to detect and replace lost qubits during syndrome extraction (SE) with minimal additional overhead. We perform circuit-level simulations to compare circuit-based [3], measurement-based [31], and Steane-based error correction [32] approaches which incorporate loss detection upon measurement, alongside other state-of-the-art approaches such as erasure conversion techniques [10, 17]. We observe that QEC performance can be substantially improved by optimizing circuit design in cases where atom loss constitutes a substantial fraction of the error budget. We then study how loss-detecting SE rounds can be optimally interleaved between transversal gates in multi-qubit deep logical Clifford circuits.
@@
We next explore an important class of circuits with minimal operations before logical measurement, which is associated with key algorithmic subroutines such as magic state distillation [33], quantum arithmetic [34], and small-angle synthesis [35, 36]. In such circuits, gate teleportation naturally detects and replaces lost qubits without additional overhead. As a result, we find that loss-detecting SE is not needed in this setting. Our numerical simulations of a teleportation-based algorithm indicate that as the frequency of gate teleportation increases, the performance is getting closer to that of erasure conversion. These results provide a comparative analysis of experimental solutions for leveraging loss and highlight that algorithmic structure plays a central role in how these factors impact performance.
@@
Before proceeding, we note that loss-to-erasure conversion has been explored at the memory level, using mid-circuit measurement in alkaline-earth-like atomic systems [17, 10, 16, 11] and superconducting qubits [19, 21, 20, 15, 37, 38, 39], as well as using leakage-reduction units with extra qubits and gates [25, 40, 41, 42]. Since the loss must be addressed in all systems and erasure conversion may not be always directly accessible after each gate operation (or, alternatively, may result in additional time and errors), we focus on the cases involving loss-resolving readout as a part of the natural QEC cycles. Most recently, atom loss-resolving readout was demonstrated experimentally and its implications for logical qubit performance were explored in Refs. [22, 24, 23, 8]. We extend these previous studies, which focus on the performance of individual logical qubits, to analyze atom loss effects at all levels ranging from decoding and QEC architecture to logical algorithms. We systematically compare different loss detection and correction methods, including their frequencies and associated space-time overheads. By focusing on logical algorithms we provide new insights into how circuit structure influences the algorithmic performance and develop new, optimized strategies for circumventing atom loss.
@@
Our paper is organized as follows. Section 2 discusses the effect of atom loss and its detection using state-selective readout, and the delayed-erasure decoder developed in this work. In Section 3, we tailor and compare different SE techniques designed to handle loss with minimal overhead, which are necessary in circuits with high depth before logical measurements. Section 4 explores the optimal frequency of interleaving loss-detecting SE with transversal gates in deep logical Clifford circuits. In Section 5, we highlight the prevalent use of teleportation in many algorithmic subroutines and evaluate its effect with and without the delayed-erasure decoder on a toy model for the small-angle synthesis algorithm. Finally, we present our conclusions in Section 6.
@@ 2 Detecting and decoding delayed erasures
Figure 1: Loss errors in logical circuits. (a) Depiction of a logical algorithm with loss-detecting SE and gate teleportation. Physical qubit losses (red crosses) can generate correlated errors within and between logical qubits, complicating error correction. (b) Space-time diagram of a logical circuit, focusing on a measure qubit lifecycle during syndrome extraction. Physical qubits progress through time, undergoing initialization, gate operations, idling, and measurement. A loss event causes future gates to be cancelled, generating correlated errors between the qubits in the gate and flipping the corresponding stabilizers.
We consider general logical algorithms during which loss is periodically detected using state-selective readout (SSR), through loss-detecting SE and loss-detecting gate teleportation, as illustrated in Fig. 1(a). SSR corresponds to directly detecting atomic qubit states |$\ket{0}$, |$\ket{1}$, or atom loss. As it occurs as part of the qubit measurement, it can be realized through various methods in neutral atoms with minimal experimental overhead [43, 44, 11, 45, 46, 47, 24, 8]. A significant challenge in addressing loss errors in comparison to erasure errors is the uncertainty in the exact error location, which complicates the decoding procedure.
As a key quantity in predicting the performance of different loss detection and correction methods, we consider the qubit lifecycle, defined as the number of circuit locations where a given qubit can potentially be lost, starting at initialization and ending at measurement. For a typical realization of two-qubit gates [14], a neutral atom qubit which was expected to interact with a lost qubit experiences a single qubit error channel, rather than an entangling gate. Therefore, losing a qubit cancels all consecutive gates, leading to correlated Clifford errors within the circuit and causing stabilizers to flip (see Fig. 1(b)). In the context of logical algorithms, this can lead to correlated errors between logical qubits due to gate cancellation between logical qubits and propagation of errors during logical entangling gates (see Fig. 1(a)).
@@
Figure 2: Delayed-erasure decoder. (a) Illustration of a qubit lifecycle and its usage in the delayed-erasure decoder. From initialization to measurement, each physical qubit can be lost at multiple possible time points, each occurring with a potentially different probability and corresponding syndrome. Upon detection of a qubit loss, the decoder accounts for these possibilities in order to improve the accuracy of the assigned correction. (b) Logical error rate for a logical memory as a function of the number of conventional SE rounds before logical measurement, here with distance $d=5$ and loss errors only with probability $p_{\text{loss}}=1\%$ per entangling gate. The delayed-erasure decoder (pink) substantially outperforms a decoder which does not account for loss information (black). Even without accurate loss information, the delayed-erasure decoder achieves comparable performance to a decoder with perfect loss information (gray, dashed).
@@
While in principle these errors can be corrected by detecting the atom loss events and treating them as erasure errors, unlike the ideal erasure channels, where the location of the loss error is precisely known, in a natural QEC cycle, the loss event is detected when the SSR is performed and can correspond to a number of different potential loss locations (Fig. 2(a)), each resulting in a potentially different set of correlated errors. These loss detection events can be viewed as delayed erasure detections, and the information from SSR can be leveraged by the appropriate decoder to improve logical performance.
@@
As a result, loss events require two adjustments to the decoder: first, when a qubit is lost, its associated stabilizer checks are no longer valid. Therefore, they need to be replaced with a so-called “supercheck”, a product of multiple stabilizer checks into a single check which is independent of the lost qubit [48, 49]. For example, consider a lost qubit $q_{4}$ that participates in neighboring stabilizers $S_{1}=Z_{1}Z_{2}Z_{3}Z_{4}$ and $S_{2}=Z_{4}Z_{5}Z_{6}Z_{7}$. The resulting supercheck operator is $S_{1}S_{2}=Z_{1}Z_{2}Z_{3}Z_{5}Z_{6}Z_{7}$, which is independent of the lost qubit $q_{4}$. Second, loss causes correlated errors, necessitating updates to the circuit error model to account for the likelihood of error propagation from the possible loss locations (see Supplementary Materials [50] Section S1 for further details).
We develop a delayed-erasure decoder which includes these two adjustments to effectively utilize imperfect information about the location of losses in time obtained from SSR. In practice, these adjustments can be paired with any decoding algorithm to augment its performance in the presence of loss. Here we primarily use the correlated most likely error (MLE) decoder from Ref. [12], as it can be applied to a broad class of codes and logical algorithms, and present results for the minimum weight perfect matching decoder in Supplementary Materials [50] Section S4.2.2 [51, 52]. Unlike previous approaches to decoding loss errors [21, 9, 22], our decoder automatically adjusts the error model based on the circuit and loss information, eliminating the need for hand-tuned models that may not easily generalize to complex logical algorithms or different SE methods.
@@
Formalizing our approach, our goal is to automatically construct a decoding graph based on the observed loss, which captures how errors (hyperedges) trigger checks (vertices that compare consecutive stabilizer measurements in time). Ideally, we would solve the MLE decoding problem, which identifies the most likely configuration of Pauli and loss errors consistent with both the observed error syndromes and loss events (see Appendix Section BB.1). However, since a full MLE solution would require considering all combinations of loss locations due to the non-additive nature of loss-induced errors, this approach is computationally intractable for even modest system sizes and low loss rates.
@@
Instead, we approximate the MLE solution by handling each loss event independently (see Appendix Section BB.2 for full details). For each loss event, we trace back the qubit’s lifecycle, accounting for all potential loss events and their associated probabilities. Each potential loss event corresponds to a loss circuit, in which certain gates are canceled due to the loss. Using Stim [53], a Clifford circuit simulator, we construct the hyperedges and their probabilities for each loss circuit. Next, we integrate all hyperedges within a qubit’s lifecycle, re-weighting them based on the probabilities of their corresponding potential loss locations. In cases of multiple losses, each loss event is calculated independently, and the results are combined to construct the final decoding graph, which serves as the input to the decoder. Empirically, this heuristic achieves performance comparable to methods that consider combinations of loss events, while significantly reducing computational overhead.
@@
To benchmark the performance of our decoder, we focus on the well-studied case of a logical memory experiment on a surface code, which involves repeated SE rounds over time, while loss is detected only by SSR. In this setting, losses on measure qubits are detected every SE round, but losses on data qubits are not detected until the logical qubit is projectively measured at the end of the circuit. We perform simulations on a single surface code with code distance $d=5$ and circuit-level noise with an entangling gate loss error of $1\%$. Unless otherwise noted, this paper considers loss and Pauli errors occurring during entangling operations, as this represents a dominant source of error in many quantum computing architectures [4, 9, 14]. Details of the error model, along with results using alternative error models, can be found in Supplementary Materials Section S4.2.3 [50] and Table 1.
We begin by analyzing the case without any loss replacement, while increasing the number of rounds—and therefore, the lifecycle length. Figure 2(b) shows the logical error rate for three types of decoders: the MLE decoder, which does not account for loss information (black lines); the delayed-erasure decoder, which uses SSR information (pink lines); and an erasure decoder with perfect loss location information (gray dashed lines). Despite lacking accurate loss information, the delayed-erasure decoder improves the logical error rate by several orders of magnitude. Furthermore, even for deep circuits with long lifecycles, its performance is comparable to that of decoding with perfect loss information. As we will see later, for shallower circuits, the performance of the delayed-erasure decoder approaches that achieved with both perfect loss location information and full lost qubit replacement every round. This is because, in shallow circuits, the propagation of loss errors is nearly deterministic, significantly simplifying the decoding problem and allowing the delayed-erasure decoder to effectively handle these errors.
In contrast, when the circuit depth is large prior to logical measurement, the qubit lifecycles are extended and loss errors can occur at multiple points. These losses introduce correlated errors, depending on when and where they occur. While the delayed-erasure decoder performs relatively well in this setting compared to a decoder which does not utilize loss information and also compared to the decoder with perfect loss information, its effectiveness deteriorates when lifecycles become too long due to error accumulation. To address this, loss detection and qubit replacement must be incorporated into the algorithm to maintain performance. Therefore, in the next section we explore practical approaches to detect and replace lost atoms with minimal overhead.

### v3

# Leveraging Qubit Loss Detection in Fault Tolerant Quantum Algorithms

@@ Abstract
Qubit loss errors constitute a dominant source of noise in many quantum hardware systems, particularly in neutral atom quantum computers. We develop a theoretical framework to effectively detect and correct loss errors in logical algorithms and leverage such loss information in decoding. Considering general quantum error correction codes and logical circuits, we introduce a delayed-erasure decoder for experimentally-motivated error models which leverages information from delayed loss detection to accurately correct loss errors, even when the precise moment of the error is unknown. Using this decoder, we identify strategies for detecting and correcting loss errors based on the logical circuit structure. For deep circuits prior to logical measurement, we explore methods to integrate loss detection into syndrome extraction with minimal overhead, identifying optimal strategies depending on the qubit loss fraction in the noise and hardware capabilities. In contrast, we find that many key algorithmic subroutines involve frequent gate teleportation, shortening the circuit depth before logical measurement and naturally replacing qubits with no additional experimental overhead. We simulate this setting using a toy model algorithm for small-angle synthesis, and find a significant performance improvement as the loss fraction increases. These results provide a path forward for advancing large-scale fault tolerant quantum computation in systems with loss error detection.
@@ 1 Introduction
Quantum error correction (QEC) is believed to be essential for realizing large-scale quantum computation, as it enables suppression of errors [1, 2, 3]. However, its practical implementation remains challenging due to its substantial resource overhead. Recent experimental advancements have demonstrated remarkable progress in implementing quantum operations across multiple logical qubits [4, 5, 6, 7, 8], and operating below error thresholds [9, 10]. These advances make it clear that practical QEC performance can be substantially improved by tailoring the error correction strategy to the particular experimental error model [11, 9, 12], choice of logic gates [4], and the structure of the algorithm itself [13, 14].
@@
In particular, qubit loss and leakage from computational subspace are dominant noise sources in many hardware systems, and can have a dramatic effect on the practical performance of quantum processors [15, 16]. In the absence of loss correction, all qubits within a QEC code can eventually disappear, destroying the encoded quantum information. Conversely, directly detecting loss provides valuable information about the error location, in contrast with Pauli errors, which are indirectly inferred from syndrome information. Recent work has shown that the rich information provided by so-called erasure detection can in fact improve QEC performance substantially [17, 18, 19, 20, 11, 12, 21, 22, 23, 24, 25, 26, 27]. Similarly, noise bias present in many platforms [15, 4, 9, 10] can also be leveraged to improve performance [28, 29, 30, 31]. These studies motivate developing strategies that incorporate loss detection and harness bias to improve practical QEC performance.
@@
In this Article, we explore the role of loss errors in error-corrected circuits across a range of quantum hardware platforms, including alkali and alkaline-earth-like neutral atoms, as well as superconducting qubits or trapped ions. We focus on the realistic scenario in which detection and correction of loss are delayed for several gate operations, and develop techniques that leverage this delayed information with minimal experimental overhead. In particular, we develop a delayed-erasure decoder to accurately interpret logical measurement results from the measured syndromes and loss detections, despite uncertainty in the exact moment of the loss error.
@@
Using this delayed-erasure decoder, we investigate the impact of loss errors in logical circuits. We demonstrate that the algorithmic structure significantly influences the optimal strategy to detect and correct loss, as summarized in Fig. 1(a). Concretely, for high-depth circuits prior to logical measurement, we develop hardware-efficient methods to detect and replace lost qubits during syndrome extraction (SE) with minimal additional overhead. We perform circuit-level simulations to compare modified circuit-based [3] and measurement-based [32] approaches which incorporate loss detection upon measurement, alongside other state-of-the-art approaches such as erasure conversion techniques [11, 18]. We observe that QEC performance can be substantially improved by optimizing circuit design to leverage loss errors in cases where it constitutes a substantial fraction of the error budget. By analyzing biased errors we find that loss has a stronger impact on performance than bias, even with bias-preserving gates. We then study how loss-detecting SE rounds can be optimally interleaved between transversal gates in multi-qubit deep logical Clifford circuits.
@@
We next find that many key algorithmic subroutines have short circuit depth before logical measurement, including magic state distillation [33], quantum arithmetic [34], and small-angle synthesis [35, 36]. In such circuits, we find that because gate teleportation naturally detects and replaces lost qubits, no loss-detecting SE is needed. We discover through numerical simulations of a teleportation-based algorithm that as the frequency of gate teleportation increases, the performance approaches that of frequent erasure conversion. These results provide a comparative analysis of experimental solutions for leveraging loss, and highlight that algorithmic structure plays a central role in how these factors impact performance.
@@
Before proceeding, we note that loss-to-erasure conversion has been explored at the memory level, using mid-circuit measurement in alkaline-earth-like atomic systems [18, 11, 17, 12] and superconducting qubits [20, 22, 21, 16, 37, 38, 39], as well as using leakage-reduction units with extra qubits and gates [26, 40, 41, 42]. More recently, loss-resolving readout was demonstrated experimentally in alkali atoms systems and dual-rail superconducting qubits and its implications for logical qubit performance were explored in Refs. [23, 25, 43, 24, 8, 10]. Here we present an end-to-end analysis of loss handling techniques, comparing all these approaches across a range of quantum computing platforms. Our study extends beyond single logical qubits to examine the role of loss at every level of computation, from decoding strategies and QEC architectures to full logical algorithms. We systematically compare different loss detection and correction methods, including their frequencies and associated space-time overheads. By focusing on logical algorithms, we provide new insights into how circuit structure influences the algorithmic performance and develop new, optimized strategies for circumventing loss errors.
@@
Our paper is organized as follows. Section 2 discusses the effect of qubit loss and its detection using state-selective readout, and the delayed-erasure decoder developed in this work. In Section 3, we tailor and compare different SE techniques designed to handle loss with minimal overhead, which are necessary in circuits with high depth before logical measurements. Section 4 explores the optimal frequency of interleaving loss-detecting SE with transversal gates in deep logical Clifford circuits. In Section 5, we highlight the prevalent use of teleportation in many algorithmic subroutines and evaluate its effect with and without the delayed-erasure decoder on a toy model for the small-angle synthesis algorithm. Finally, we present our conclusions in Section 6.
@@ 2 Detecting and decoding delayed erasures
Figure 1: Loss errors in logical circuits. (a) Depiction of a logical algorithm with loss-detecting SE and gate teleportation. Physical qubit losses (red crosses) can generate correlated errors within and between logical qubits. (b) Space-time diagram of a logical circuit, focusing on a measure qubit lifecycle during syndrome extraction. Physical qubits progress through time, undergoing initialization, gate operations, idling, and measurement. A loss event causes future gates to be canceled, generating correlated errors between the qubits in the gate and flipping the corresponding stabilizers.
[…]
We explicitly differentiate between two closely related error types. We define a loss error as one in which a qubit leaves the computational subspace and becomes unavailable until detection and replacement, encompassing both loss and leakage (depending on hardware platform). An erasure error, by contrast, is a loss that is detected at the gate layer where it occurs and is immediately reinitialized; both the time and qubit index are known to the decoder.
Loss events lead to correlated errors: in neutral-atom systems, a lost qubit cancels subsequent gates, while its intended interaction partner may experience an effective single-qubit error [15]; in superconducting platforms, a qubit interacting with a lost qubit suffers from depolarizing noise [21]. In both cases, correlated errors can propagate within or across logical qubits during entangling operations (see Fig. 1(a,b)), and must be handled carefully by the decoder. Here we adopt the gate-cancellation model for loss.
@@
Figure 2: Delayed-erasure decoder. (a) Illustration of a qubit lifecycle and its usage in the delayed-erasure decoder. From initialization to measurement, each physical qubit can be lost at multiple possible time points, each occurring with a potentially different probability and corresponding syndrome. Upon detection of a qubit loss, the decoder accounts for these possibilities in order to improve the accuracy of the assigned correction. (b) Logical error rate for a logical memory as a function of the number of conventional SE rounds before logical measurement, here with distance $d=5$ and loss errors with probability $p_{\text{loss}}=1\%$ per entangling gate. The delayed-erasure decoder (pink) substantially outperforms a decoder which does not account for loss information (black). Even without loss moment information, the delayed-erasure decoder achieves comparable performance to a decoder with perfect loss time-location (gray, dashed).
@@
While in principle these errors can be corrected by detecting and replacing the lost qubits and treating them as erasure errors, unlike the ideal erasure channels, where the moment in time of the loss error is precisely known, in a natural QEC cycle, the loss event is detected when the SSR or other loss detection mechanism is performed and can correspond to a number of different potential loss locations (Fig. 2(a)), each resulting in a potentially different set of correlated errors. These loss detection events can be viewed as delayed erasure detections, and the information revealed by the detection mechanism can be leveraged by the appropriate decoder to improve logical performance.
@@
As a result, loss events require two adjustments to the decoder: first, when a qubit is lost, its associated stabilizer checks are no longer valid. Therefore, they need to be replaced with a so-called “supercheck”, a product of multiple stabilizer checks into a single check which is independent of the lost qubit [51, 52]. For example, consider a lost qubit $q_{4}$ that participates in neighboring stabilizers $S_{1}=Z_{1}Z_{2}Z_{3}Z_{4}$ and $S_{2}=Z_{4}Z_{5}Z_{6}Z_{7}$. The resulting supercheck operator is $S_{1}S_{2}=Z_{1}Z_{2}Z_{3}Z_{5}Z_{6}Z_{7}$, which is independent of the lost qubit $q_{4}$. Second, loss causes correlated errors, necessitating updates to the circuit error model to account for the likelihood of error propagation from the possible loss locations (see Appendix Section A).
@@
We develop a delayed-erasure decoder which includes these two adjustments to effectively utilize imperfect information about the location of losses in time obtained from SSR. As a key quantity in predicting the performance of different loss detection and correction methods, we consider the qubit lifecycle, defined as the number of circuit locations where a given qubit can potentially be lost, starting at initialization and ending at measurement. Formalizing our approach, our goal is to automatically construct a decoding hypergraph based on the observed loss, which captures how errors (hyperedges) trigger checks (vertices that compare consecutive stabilizer checks in time) and is the input to the decoder. Ideally, we would solve the most-likely-error (MLE) decoding problem, which identifies the most likely configuration of Pauli and loss errors consistent with both the observed error syndromes and loss events (see Appendix Section B.B.1). However, a full MLE solution would require considering all combinations of loss locations, since loss-induced errors can produce non-Pauli, correlated error patterns, making them non-additive, as observed numerically.
@@
Instead, we approximate the MLE solution by handling each loss event independently (see Appendix Section BB.2 for full details). For each loss event, we trace back the qubit’s lifecycle, accounting for all potential loss events and their associated probabilities. Each potential loss event corresponds to a loss circuit, in which certain gates are canceled due to the loss. Using Stim [53], a Clifford circuit simulator, we construct the hyperedges and their probabilities for each loss circuit. Next, we integrate all hyperedges within a qubit’s lifecycle, re-weighting them based on the probabilities of their corresponding potential loss locations. In cases of multiple losses, each loss event is calculated independently, and the results are combined to construct the final decoding hypergraph used by the decoder, calculated by:
$\sum_{i}{D}_{i}+{D}_{\mathrm{Pauli}}+\omega\cdot{D}_{\mathrm{first\ comb}},$ (1)
where $D_{i}$ is the decoding hypergraph for a lossy lifecycle $i$, $D_{\mathrm{Pauli}}$ is the decoding hypergraph for Pauli errors generated from the lossless circuit, and $D_{\mathrm{first\ comb}}$ is the decoding hypergraph built from the earliest potential loss in each lossy lifecycle $i$. The weighting parameter $\omega$ tunes the contribution of $D_{\mathrm{first\ comb}}$, and is set to $\omega=0$ in the main-text simulations (see Appendix B.2 for details). Empirically, this heuristic achieves performance comparable to methods that consider combinations of loss events, while significantly reducing computational overhead.
The updated decoding hypergraph can then be processed using a standard decoder appropriate for the QEC code in use. For example, one may use a MLE decoder, minimum weight perfect matching (MWPM) for topological codes [49, 54, 9], belief propagation with ordered statistics decoding (BP+OSD) for sparse codes [55], or machine learning-based decoders [56, 9]. Unlike previous approaches to decoding loss errors [22, 9, 23], our decoder automatically adjusts the error model based on any general circuit and loss information, eliminating the need for hand-tuned models that may not easily generalize to complex logical algorithms or different SE methods. Here we primarily use the MLE decoder from Ref. [13], as it can decode generic stabilizer codes and logical algorithms. We also present results for the MWPM decoder in Appendix Section C.
@@
To benchmark the decoder, we use a surface-code logical memory with repeated SE rounds in which loss is detected only by SSR: losses on measure qubits are detected each round, whereas losses on data qubits are detected only at the final projective readout. Unless otherwise noted, this paper considers loss and Pauli errors occurring during entangling operations, as this represents a dominant source of error in many quantum computing architectures [4, 9, 15]. Details of the error model, along with results using alternative error models, can be found in Appendix F and Table 1. As shown in Fig. 2(b), a delayed-erasure decoder that uses only SSR information lowers the logical error rate by orders of magnitude relative to an MLE decoder that ignores loss information and remains comparable to a decoder with perfect loss time-location, even for deeper circuits with longer lifecycles.